# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ 4.2 Quadratic regression bound
As noted in Section 3.1, the quadratic regression can be converted to linear regression using $\|h\|_{P}^{2}=\left\langle hh^{\top},P\right\rangle_{F}=\left\langle\mathrm{svec}(hh^{\top}),\mathrm{svec}(P)\right\rangle$. To analyze this linear regression with an intercept, we need the following lemma. We note that a similar lemma without considering the intercept has been proved in (Jadbabaie et al., 2021, Proposition).
@@
Let $(h_{0}^{(i)})_{i=1}^{n}$ be $n$ independent observations of the $d$-dimensional random vector $h_{0}\sim\mathcal{N}(0,I_{d})$. Let $f_{0}^{(i)}:=\mathrm{svec}(h_{0}^{(i)}(h_{0}^{(i)})^{\top})$ and $\overline{f}_{0}^{(i)}:=[f_{0}^{(i)};1]$. There exists an absolute constant $a>0$, such that as long as $n\geq ar^{4}\log(ar^{2}/p)$, with probability at least $1-p$,
@@
Let $f_{0}=\mathrm{svec}(h_{0}h_{0}^{\top})$ and $\overline{f}_{0}=[f_{0};1]$. We first show that $\lambda_{\min}(\mathbb{E}[\overline{f}\;\overline{f}^{\top}])$ is lower bounded. Consider
$\overline{\Sigma}=\mathbb{E}[\overline{f}\;\overline{f}^{\top}]=\begin{bmatrix}2I_{d(d-1)/2}&0&0\\ 0&2I_{d}+1_{d}1_{d}^{\top}&1_{d}\\ 0&1_{d}^{\top}&1\end{bmatrix}.$
@@
Then, with the similar concentration arguments to those in (Jadbabaie et al., 2021, Appendix A.1), we can show that with probability at least $1-p$,
$\lambda_{\min}(F_{0}^{\top}F_{0})=\Omega(d^{-1}n).$
Lemma 1 lower bounds the minimum singular value of a matrix that contains the fourth powers of elements in standard Gaussian random vectors. The following lemma is the main result in Section 4.2.
@@
Define random variable $c:=(h^{\ast})^{\top}N^{\ast}h^{\ast}+b^{\ast}+e$, where $h^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$ is a $d$-dimensional Gaussian random vector, $N^{\ast}\in\mathbb{R}^{d\times d}$ is a positive semidefinite matrix, $b^{\ast}\in\mathbb{R}$ is a constant and $e$ is a zero-mean subexponential random variable with $\|e\|_{\psi_{1}}\leq E$. Assume that $\|N^{\ast}\|_{2}$ and $\|\Sigma_{\ast}\|_{2}$ are $\mathcal{O}(1)$. Let $\sigma_{\min}(\Sigma_{\ast}^{1/2})\geq\beta>0$. Assume that $\beta=\Omega(1)$. Define $h:=h^{\ast}+\delta$ where the noise vector $\delta$ can be correlated with $h^{\ast}$ and its $\ell_{2}$ norm is sub-Gaussian with $\mathbb{E}[\|\delta\|]\leq\epsilon$, $\|\|\delta\|\|_{\psi_{2}}\leq\epsilon$. Assume that $\epsilon\leq\min((d\|\Sigma_{\ast}\|^{1/2}),a(\beta\wedge 1)d^{-3/2}\|\Sigma_{\ast}\|_{2}^{-1/2}(\log(n/p))^{-1})$ for some absolute constant $a>0$. Suppose we get $n$ observations $h^{(i)}$ and $c^{(i)}$ of $h$ and $c$, where $((h^{\ast})^{(i)})_{i=1}^{n}$ are independent and $(\delta^{(i)})_{i=1}^{n}$ can be correlated. Consider the regression problem
@@
There exists an absolute constant $a_{0}>0$, such that as long as $n\geq a_{0}d^{4}\log(a_{0}d^{2}/p)\log(1/p)$, with probability at least $1-p$, $\|\hat{N}-N^{\ast}\|_{F}$ and $|\hat{b}-b^{\ast}|$ are bounded by
$\mathcal{O}((d(1+Ed^{1/2})\log(n/p)+\|b^{\ast}\|)\epsilon+d^{2}En^{-1/2}\log^{1/2}(1/p)).$
@@
Let $f^{(i)}:=\mathrm{svec}(h^{(i)}(h^{(i)})^{\top})$ denote the covariates and $\overline{f}^{(i)}=[f^{(i)};1]$ denote the extended covariates. Define $(f^{\ast})^{(i)}$ and $(\overline{f}^{\ast})^{(i)}$ similarly by replacing $h^{(i)}$ with $(h^{\ast})^{(i)}$. Then, regression (4.3) can be written as
@@
Let $\overline{F}:=[\overline{f}^{(1)},\ldots,\overline{f}^{(n)}]^{\top}$ be an $n\times\frac{d(d+3)}{2}$ matrix whose $i$th row is $(\overline{f}^{(i)})^{\top}$. Define $\overline{F}^{\ast}$ similarly by replacing $\overline{f}^{(i)}$ with $(\overline{f}^{\ast})^{(i)}$. Solving linear regression (4.4) gives
@@
$\sigma_{\min}(\Phi^{\ast})=\sigma_{\min}(\Sigma_{\ast}^{1/2})^{2}=\sigma_{\min}(\Sigma_{\ast})=\beta>0.$
By Lemma 1, there exist absolute constants $a_{0},a_{1}>0$, such that if $n\geq a_{0}d^{4}\log(a_{0}d^{2}/p)$, with probability at least $1-p$, $\lambda_{\min}(\overline{F}_{0}^{\top}\overline{F}_{0})\geq a_{1}d^{-1}n$. Since $\overline{f}=\textnormal{diag}(\Phi^{\ast},1)\overline{f}_{0}$ and $\overline{F}^{\ast}=\overline{F}_{0}\textnormal{diag}((\Phi^{\ast})^{\top},1)$,
@@
By Weyl’s inequality,
@@
=\;= $\|h^{\ast}(h^{\ast}-h)^{\top}+(h^{\ast}-h)h^{\top}\|_{2}$
[…]
Since $h^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$, $\|h^{\ast}\|$ is sub-Gaussian with its mean and sub-Gaussian norm bounded by $\mathcal{O}((d\|\Sigma_{\ast}\|)^{1/2})$. Since $\|\delta\|$ is sub-Gaussian with its mean and sub-Gaussian norm bounded by $\epsilon\leq(d\|\Sigma_{\ast}\|)^{1/2}$, we conclude that $\|h^{\ast}(h^{\ast})^{\top}-hh^{\top}\|_{F}$ is subexponential with its mean and subexponential norm bounded by $\mathcal{O}(\epsilon(d\|\Sigma_{\ast}\|)^{1/2})$. Hence, with probability at least $1-p$,
$\|h^{\ast}(h^{\ast})^{\top}-hh^{\top}\|_{F}^{2}=\mathcal{O}(\epsilon^{2}d\|\Sigma_{\ast}\|\log^{2}(n/p)).$
Therefore,
@@
[…]
@@
$\sigma_{\min}(\overline{F})^{-1}\|(\overline{F}^{\ast}-\overline{F})[\mathrm{svec}(N^{\ast});b^{\ast}]\|=\;$ $\mathcal{O}(\sigma_{\min}(\overline{F})^{-1})\|(F^{\ast}-F)\mathrm{svec}(N^{\ast})+1_{n}b^{\ast}]\|$
=\;= $\mathcal{O}((d^{1/2}n^{-1/2})(\|(F^{\ast}-F)\mathrm{svec}(N^{\ast})\|+\|1_{n}b^{\ast}\|).$
Using arguments similar to those in (Mhammedi et al., 2020, Section B.2.13), we have
$\|(F^{\ast}-F)\mathrm{svec}(N^{\ast})\|^{2}=\;$ $\sum_{i=1}^{n}\left\langle\mathrm{svec}((h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top})-\mathrm{svec}(h^{(i)}(h^{(i)})^{\top}),\mathrm{svec}(N^{\ast})\right\rangle$
=$\overset{(i)}{=}\;$ $\sum\nolimits_{i=1}^{n}\left\langle(h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top}-h^{(i)}(h^{(i)})^{\top},N^{\ast}\right\rangle_{F}$
@@
=\;= $2\|N^{\ast}\|_{2}^{2}\|F^{\ast}-F\|_{F}^{2},$
where $\left\langle\cdot,\cdot\right\rangle_{F}$ denotes the Frobenius product between matrices in $(i)$, $\|\cdot\|_{\ast}$ denotes the nuclear norm in $(ii)$, and $(iii)$ follows from the fact that the matrix $(h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top}-h^{(i)}(h^{(i)})^{\top}$ has at most rank two. Combining with $\|1_{n}b^{\ast}\|=n^{1/2}\|b^{\ast}\|$, we obtain that the term $(a)$ in (4.7) is bounded by
$\mathcal{O}\big{(}d^{1/2}n^{-1/2}\cdot(\|N^{\ast}\|_{2}\cdot\epsilon(d\|\Sigma_{\ast}\|_{2})^{1/2}n^{1/2}\log(n/p)+n^{1/2}\|b^{\ast}\|)\big{)}$
=\;= $\mathcal{O}(d\log(n/p)+\|b^{\ast}\|)\epsilon).$
@@
$(b)=\|\overline{F}^{\dagger}\xi\|\leq\sigma_{\min}(\overline{F}^{\top}\overline{F})^{-1}\|\overline{F}^{\top}\xi\|=\mathcal{O}(dn^{-1})(\|(\overline{F}-\overline{F}^{\ast})^{\top}\xi\|+\|(\overline{F}^{\ast})^{\top}\xi\|).$
@@
=\;= $\mathcal{O}(\epsilon(d\|\Sigma_{\ast}\|_{2}n)^{1/2}\log(n/p))\cdot\mathcal{O}(En^{1/2}\log(n/p))$
=\;= $\mathcal{O}(d^{1/2}En\log^{2}(n/p)\epsilon).$
To bound $\|(\overline{F}^{\ast})^{\top}\xi\|=\|\textnormal{diag}(\Phi^{\ast},1)\overline{F}_{0}^{\top}\xi\|$, note that $\|\overline{F}_{0}^{\top}\xi\|=\|\sum_{i=1}^{n}\overline{f}_{0}^{(i)}e^{(i)}\|$. Consider the $j$th component in the summation $\sum_{i=1}^{n}[\overline{f}_{0}^{(i)}]_{j}(e^{\prime})^{(i)}$. Recall that $f_{0}=\mathrm{svec}(h_{0}h_{0}^{\top})$, so $[\overline{f}_{0}]_{j}$ is either the square of a standard Gaussian random variable, $\sqrt{2}$ times the product of two independent standard Gaussian random variables, or one. Hence, $[f_{0}]_{j}$ is subexponential with mean and $\|[f_{0}]_{j}\|_{\psi_{1}}$ both bounded by $\mathcal{O}(1)$. As a result, the product $[f_{0}]_{j}e$ is $\frac{1}{2}$-sub-Weibull, with the sub-Weibull norm being $\mathcal{O}(E)$. By (Hao et al., 2019, Theorem 3.1),
@@
Hence, the norm of the $d(d+1)/2$-dimensional vector $F_{0}^{\top}\xi$ is $\mathcal{O}(dEn^{1/2}\log^{1/2}(1/p))$. By the properties of the symmetric Kronecker product (Schacke, 2004), $\|\Phi^{\ast}\|_{2}=\|\Sigma_{\ast}^{1/2}\|_{2}^{2}=\|\Sigma_{\ast}\|_{2}=\mathcal{O}(1)$. Then,
$\|(\overline{F}^{\ast})^{\top}\xi\|=\mathcal{O}(dEn^{1/2}\log^{1/2}(1/p)).$
@@
$\mathcal{O}(dn^{-1})\cdot\mathcal{O}(d^{1/2}En\log(n/p)\epsilon+dEn^{1/2}\log^{1/2}(1/p))$
=\;= $\mathcal{O}\big{(}d^{3/2}E\log(n/p)\epsilon+d^{2}En^{-1/2}\log^{1/2}(1/p)\big{)}.$
@@
=\;= $\mathcal{O}((d\log(n/p)+\|b^{\ast}\|+d^{3/2}E\log(n/p))\epsilon+d^{2}En^{-1/2}\log^{1/2}(1/p))$
=\;= $\mathcal{O}((d(1+Ed^{1/2})\log(n/p)+\|b^{\ast}\|)\epsilon+d^{2}En^{-1/2}\log^{1/2}(1/p)),$

### v3

@@ 4.2 Quadratic regression bound
As noted in §3.1, the quadratic regression can be converted to linear regression using $\|h\|_{P}^{2}=\left\langle hh^{\top},P\right\rangle_{F}=\left\langle\mathrm{svec}(hh^{\top}),\mathrm{svec}(P)\right\rangle$. To analyze this linear regression with an intercept, we need the following lemma. We note that a similar lemma without considering the intercept has been proved in Jadbabaie et al. (2021, Proposition 1).
@@
Let $(h_{0}^{(i)})_{i=1}^{n}$ be $n$ independent observations of the $d$-dimensional random vector $h_{0}\sim\mathcal{N}(0,I_{d})$. Let $f_{0}^{(i)}:=\mathrm{svec}(h_{0}^{(i)}(h_{0}^{(i)})^{\top})$ and $\overline{f}_{0}^{(i)}:=[f_{0}^{(i)};1]$. There exists an absolute constant $a>0$, such that as long as $n\geq ar^{4}\log(ar^{2}/p)$, with probability at least 1p1-p,
@@
Let $f_{0}=\mathrm{svec}(h_{0}h_{0}^{\top})$ and $\overline{f}_{0}=[f_{0};1]$. We first show that $\lambda_{\min}(\mathbb{E}[\overline{f}_{0}\overline{f}_{0}^{\top}])$ is lower bounded. Consider
$\overline{\Sigma}=\mathbb{E}[\overline{f}_{0}\overline{f}_{0}^{\top}]=\begin{bmatrix}2I_{d(d-1)/2}&0&0\\ 0&2I_{d}+1_{d}1_{d}^{\top}&1_{d}\\ 0&1_{d}^{\top}&1\end{bmatrix}$.
@@
Then, with the similar concentration arguments to those in Jadbabaie et al. (2021, Appendix A.1), we can show that with probability at least 1p1-p,
$\lambda_{\min}(F_{0}^{\top}F_{0})=\Omega(d^{-1}n),$
which completes the proof. ∎
Lemma 1 lower bounds the minimum singular value of a matrix that contains the fourth powers of elements in standard Gaussian random vectors. The following lemma is the main result in § 4.2.
@@
Define random variable $c:=(h^{\ast})^{\top}N^{\ast}h^{\ast}+b^{\ast}+e$, where $h^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast})$ is a $d$-dimensional Gaussian random vector, $N^{\ast}\in\mathbb{R}^{d\times d}$ is a positive semidefinite matrix, $b^{\ast}\in\mathbb{R}$ is a constant and $e$ is a zero-mean subexponential random variable with $\|e\|_{\psi_{1}}\leq E$. Assume that $\|N^{\ast}\|_{2}=\mathcal{O}(1)$ and that $\lambda_{\min}(\Sigma_{\ast}^{1/2})\geq\beta=\Omega(1)$. Define $h:=h^{\ast}+\delta$ where the perturbation vector $\delta$ can be correlated with $h^{\ast}$ and its $\ell_{2}$ norm is sub-Gaussian with $\mathbb{E}[\|\delta\|]\leq\epsilon$, $\|\|\delta\|\|_{\psi_{2}}\leq\epsilon$. Assume that $\epsilon\leq\min((d\|\Sigma_{\ast}\|_{2})^{1/2},a(\beta\wedge 1)d^{-3/2}\|\Sigma_{\ast}\|_{2}^{-1/2}(\log(n/p))^{-1})$ for some absolute constant $a>0$. Suppose we get $n$ observations $h^{(i)}$ and $c^{(i)}$ of $h$ and $c$, where $((h^{\ast})^{(i)})_{i=1}^{n}$ are independent and $(\delta^{(i)})_{i=1}^{n}$ can be correlated. Consider the regression problem
@@
There exists an absolute constant $a_{0}>0$, such that as long as $n\geq a_{0}d^{4}\log(a_{0}d^{2}/p)\log(1/p)$, with probability at least 1p1-p, $\|\hat{N}-N^{\ast}\|_{F}$ and $|\hat{b}-b^{\ast}|$ are bounded by
$\mathcal{O}\big{(}\epsilon\|\Sigma_{\ast}\|_{2}^{1/2}(d\log(n/p)+d^{3/2}E\log^{2}(n/p))+d^{2}\|\Sigma_{\ast}\|_{2}En^{-1/2}\log^{1/2}(1/p)\big{)}.$
@@
Let $f^{(i)}:=\mathrm{svec}(h^{(i)}(h^{(i)})^{\top})$ denote the covariates and $\overline{f}^{(i)}:=[f^{(i)};1]$ denote the extended covariates. Define $(f^{\ast})^{(i)}$ and $(\overline{f}^{\ast})^{(i)}$ similarly by replacing $h^{(i)}$ with $(h^{\ast})^{(i)}$. Then, regression (4.1) can be written as
@@
Let $\overline{F}:=[\overline{f}^{(1)},\ldots,\overline{f}^{(n)}]^{\top}$ be the $n\times\frac{d(d+3)}{2}$ matrix whose $i$th row is $(\overline{f}^{(i)})^{\top}$. Define $\overline{F}^{\ast}$ similarly by replacing $\overline{f}^{(i)}$ with $(\overline{f}^{\ast})^{(i)}$. Solving linear regression (4.2) gives
@@
$\sigma_{\min}(\Phi^{\ast})=\lambda_{\min}(\Sigma_{\ast}^{1/2})^{2}=\lambda_{\min}(\Sigma_{\ast})=\beta>0.$
By Lemma 1, there exist absolute constants $a_{0},a_{1}>0$, such that if $n\geq a_{0}d^{4}\log(a_{0}d^{2}/p)$, with probability at least 1p1-p, $\lambda_{\min}(\overline{F}_{0}^{\top}\overline{F}_{0})\geq a_{1}d^{-1}n$. Since $\overline{f}=\textnormal{diag}(\Phi^{\ast},1)\overline{f}_{0}$ and $\overline{F}^{\ast}=\overline{F}_{0}\textnormal{diag}((\Phi^{\ast})^{\top},1)$,
@@
By Weyl’s inequality for singular values,
@@
$=\;$ $\sqrt{2}\|h^{\ast}(h^{\ast}-h)^{\top}+(h^{\ast}-h)h^{\top}\|_{2}$
$\leq\;$ $\sqrt{2}(\|h^{\ast}\|+\|h\|)\|\delta\|.$
Since $^{\ast}\sim\mathcal{N}(0,\Sigma_{\ast}), h\|h^{\ast}\|$ is sub-Gaussian with its mean and sub-Gaussian norm bounded by $\mathcal{O}((d\|\Sigma_{\ast}\|)^{1/2})$. Since $\|\delta\|$ is sub-Gaussian with its mean and sub-Gaussian norm bounded by $\epsilon\leq(d\|\Sigma_{\ast}\|)^{1/2}$, we conclude that $\|h^{\ast}(h^{\ast})^{\top}-hh^{\top}\|_{F}$ is subexponential with its mean and subexponential norm bounded by $\mathcal{O}(\epsilon(d\|\Sigma_{\ast}\|)^{1/2})$. Hence, with probability at least 1p1-p,
$\|h^{\ast}(h^{\ast})^{\top}-hh^{\top}\|_{F}^{2}=\mathcal{O}(\epsilon^{2}d\|\Sigma_{\ast}\|_{2}\log^{2}(1/p)).$
Therefore, by the union bound over $n$ observations,
@@
$\leq$ $\underbrace{\|\overline{F}^{\dagger}(\overline{F}^{\ast}-\overline{F})[\mathrm{svec}(N^{\ast});b^{\ast}]\|}_{(a)}+\underbrace{\|\overline{F}^{\dagger}\xi\|}_{(b)}.$
@@
$\sigma_{\min}(\overline{F})^{-1}\|(\overline{F}^{\ast}-\overline{F})[\mathrm{svec}(N^{\ast});b^{\ast}]\|=\;$ $\mathcal{O}(\sigma_{\min}(\overline{F})^{-1})\|(F^{\ast}-F)\mathrm{svec}(N^{\ast})\|$
$=\;$ $\mathcal{O}((d^{1/2}n^{-1/2})\|(F^{\ast}-F)\mathrm{svec}(N^{\ast})\|.$
Using arguments similar to those in Mhammedi et al. (2020, Section B.2.13), we have
$\|(F^{\ast}-F)\mathrm{svec}(N^{\ast})\|^{2}$
$=\;$ $\sum_{i=1}^{n}\left\langle\mathrm{svec}((h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top})-\mathrm{svec}(h^{(i)}(h^{(i)})^{\top}),\mathrm{svec}(N^{\ast})\right\rangle^{2}$
$\overset{(i)}{=}\;$ $\sum\nolimits_{i=1}^{n}\left\langle(h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top}-h^{(i)}(h^{(i)})^{\top},N^{\ast}\right\rangle_{F}^{2}$
@@
$=\;$ $2\|N^{\ast}\|_{2}^{2}\|F^{\ast}-F\|_{F}^{2},$
where $\left\langle\cdot,\cdot\right\rangle_{F}$ denotes the Frobenius product between matrices in $(i)$, $\|\cdot\|_{\ast}$ denotes the nuclear norm in $(ii)$, and $(iii)$ follows from the fact that the matrix $(h^{\ast})^{(i)}((h^{\ast})^{(i)})^{\top}-h^{(i)}(h^{(i)})^{\top}$ has at most rank two. Hence, term $(a)$ in (4.5) is bounded by
$\mathcal{O}\big{(}d^{1/2}n^{-1/2}\|N^{\ast}\|_{2}\cdot\epsilon(d\|\Sigma_{\ast}\|_{2})^{1/2}n^{1/2}\log(n/p)\big{)}=\mathcal{O}\big{(}\epsilon d\|\Sigma_{\ast}\|_{2}^{1/2}\log(n/p)\big{)}.$
@@
$(b)=\|\overline{F}^{\dagger}\xi\|\leq\lambda_{\min}(\overline{F}^{\top}\overline{F})^{-1}\|\overline{F}^{\top}\xi\|=\mathcal{O}(dn^{-1})(\|(\overline{F}-\overline{F}^{\ast})^{\top}\xi\|+\|(\overline{F}^{\ast})^{\top}\xi\|).$
@@
$=\;$ $\mathcal{O}(\epsilon(d\|\Sigma_{\ast}\|_{2}n)^{1/2}\log(n/p))\cdot\mathcal{O}(En^{1/2}\log(n/p))$
$=\;$ $\mathcal{O}(\epsilon(d\|\Sigma_{\ast}\|_{2})^{1/2}En\log^{2}(n/p)).$
To bound $\|(\overline{F}^{\ast})^{\top}\xi\|=\|\textnormal{diag}(\Phi^{\ast},1)\overline{F}_{0}^{\top}\xi\|$, note that $\|\overline{F}_{0}^{\top}\xi\|=\|\sum_{i=1}^{n}\overline{f}_{0}^{(i)}e^{(i)}\|$. Consider the $j$th component in the summation $\sum_{i=1}^{n}[\overline{f}_{0}^{(i)}]_{j}(e^{\prime})^{(i)}$. Recall that $f_{0}=\mathrm{svec}(h_{0}h_{0}^{\top})$, so $\overline{f}_{0}]_{j}$ is either the square of a standard Gaussian random variable, $\sqrt{2}$ times the product of two independent standard Gaussian random variables, or one. Hence, $f_{0}]_{j}$ is subexponential with mean and $\|[f_{0}]_{j}\|_{\psi_{1}}$ both bounded by $\mathcal{O}(1)$. As a result, the product $f_{0}]_{j}e$ is $\frac{1}{2}$-sub-Weibull, with the sub-Weibull norm being $\mathcal{O}(E)$. By Hao et al. (2019, Theorem 3.1),
@@
Hence, the norm of the $d(d+1)/2$-dimensional vector $F_{0}^{\top}\xi$ is $\mathcal{O}(dEn^{1/2}\log^{1/2}(1/p))$. By the properties of the symmetric Kronecker product (Schacke, 2004), $\|\Phi^{\ast}\|_{2}=\|\Sigma_{\ast}^{1/2}\|_{2}^{2}=\|\Sigma_{\ast}\|_{2}$. Then,
$\|(\overline{F}^{\ast})^{\top}\xi\|=\mathcal{O}(d\|\Sigma_{\ast}\|_{2}En^{1/2}\log^{1/2}(1/p)).$
@@
$\mathcal{O}(dn^{-1})\cdot\mathcal{O}((d\|\Sigma_{\ast}\|_{2})^{1/2}En\log^{2}(n/p)\epsilon+dn^{1/2}\|\Sigma_{\ast}\|_{2}E\log^{1/2}(1/p))$
$=\;$ $\mathcal{O}\big{(}\epsilon d^{3/2}\|\Sigma_{\ast}\|_{2}^{1/2}E\log^{2}(n/p)+d^{2}\|\Sigma_{\ast}\|_{2}En^{-1/2}\log^{1/2}(1/p)\big{)}.$
@@
$=\;$ $\mathcal{O}\big{(}\epsilon\|\Sigma_{\ast}\|_{2}^{1/2}(d\log(n/p)+d^{3/2}E\log^{2}(n/p))+d^{2}\|\Sigma_{\ast}\|_{2}En^{-1/2}\log^{1/2}(1/p)\big{)},$